OpenStudy (anonymous):

Asanji can clean an attic in 13 hrs. jenny can clean the same attic in 16 hrs. if they worked together,how long would it take them?

OpenStudy (anonymous):

dumbcow can you help me out with another question of mine again its related to the one i asked you before after your finished :) heres the link: http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4dda06180ada8b0b363243c7

OpenStudy (dumbcow):

shortcut is (16)(13)/(16+13) =208/29 =7.17 hours

OpenStudy (anonymous):

The two can clean together in 208/29=7.17 hrs(approx). Reason: Asanji can complete 1/13 in one hr. Jenny can complete 1/16 of the work in one hr. Both can complete (1/13)+(1/16)=29/208 part in one hr. Therefore both can complete the whole work in 208/29=7.17 hrs.